OpenStudy (anonymous):

If f(x)=-x+5 and g(x)=x^2, what is (g of f)(-1)?

OpenStudy (anonymous):

(g o f)(-1) = g(f(-1)) In other words, plug -1 into your f function then take this result and plug it into g. f(-1) = -(-1) + 5 = 1 + 5 = 6 So now you take this and plug it into g...so you need g(6) = 6^2 = ...

OpenStudy (anonymous):

Can you simplify?

OpenStudy (lalaly):

f(-1)=-(-1)+5 = 6 g(6)= 6^2 = 36

OpenStudy (lalaly):

f(g(-1)=36

OpenStudy (anonymous):

I can't get any more simple, without giving you the answer. All that is left is calculating 6^2...which lalaly went ahead and did for you.